# Understanding Oversmoothing in GNNs as Consensus in Opinion Dynamics

### v1

@@ Appendix B Datasets
Experiments in Section 6 performs semi-supervised node classification task on the Cora (McCallum et al., 2000), Citeseer (Sen et al., 2008), Pubmed (Namata et al., 2012), CoauthorCS (Shchur et al., 2018) and Amazon Computers and Photo (McAuley et al., 2015) datasets implemented by PyTorch Geometric (Fey & Lenssen, 2019). The statistics of these datasets are summarized in Table 4. Experiments in Section 6 implements the full dataset while recent continuous-depth models such as $-\ell$ and KuramotoGNN makes use of the largest connected component of the dataset.
Table 4: Statistics of datasets. A summary of key characteristics for the 6 datasets presented in Section 6.2.
Dataset Graph Classes Features Nodes Edges
Cora 1 7 1433 2708 5278
Citeseer 1 6 3703 3312 4552
Pubmed 1 3 500 19717 44324
CoauthorCS 1 15 6805 18333 81894
Computers 1 10 767 13381 245778
Photo 1 8 745 7487 119043
B.1 Cora
@@
The Cora dataset contains a citation graph where 2708 computer science publications are connected by 5278 citation edges. Each publication has an 1433-dimensional bag-of-words vector derived from a paper keyword dictionary. Publications are classified into one of 7 classes corresponding to their primary research area.
B.2 Citeseer
@@
The Citeseer dataset contains a citation graph where 3312 computer science publications are connected by 4552 citation edges. Each publication has a 3703-dimensional bag-of-words vector derived from a paper keyword dictionary. Publications are classified into one of 6 classes corresponding to their primary research area.
B.3 Pubmed
@@
The Pubmed dataset contains a ciation graph where 19717 biomedical publications are connected by 44324 citation edges. Each publication is represented by a 500-dimensional TF/IDF weighted word vector derived from a paper keyword dictionary. Publications are classified into one of 3 classes corresponding to their primary research area.
B.4 CoauthorCS
@@
The CoauthorCS dataset is one segment of the Coauthor Graph datasets that contains a co-authorship graph that consist of 18333 authors and connected by 81894 co-authorship edges. Each author is represented by a 6805-dimensional bag-of-words feature vector derived from their paper keywords. Authors are classified into one of 15 classes corresponding to their primary research area.
B.5 Amazon Computers
@@
The Amazon Computers dataset, denoted as Computers in our paper, contains a co-purchase graph where 13381 computer products are connected by 81894 edges. The edges indicate that two products are frequently bought. Each product is represented by a 767-dimensional bag-of-words feature vector derived from their product reviews. Products are classified into one of 10 classes corresponding to their product categories.
B.6 Amazon Photo
@@
The Photo dataset, denoted as Photo in our paper, contains a co-purchase graph where 7487 photo products are connected by 119043 edges. The edges indicate that two products are frequently bought. Each product is represented by a 745-dimensional bag-of-words feature vector derived from their product reviews. Products are classified into one of 8 classes corresponding to their product categories.
Appendix C Experimental Details
C.1 Structure of BIMP
Our model BIMP comprises a learnable encoder $\phi$ and a learnable decoder $\psi$, connected by a learnable opinion exchange process.
The encoder, $\phi:\mathbb{R}^{d_{\text{in}}}\to\mathbb{R}^{N_{o}}$, is represented by a fully connected neural network that maps node features of dimension $d_{\text{in}}$ into opinions of dimension $N_{o}$. The initial embedding, serving as the initial state $\mathbf{X}(0)$, is input into an evolutionary ODE
[…]
where $\frac{d\mathbf{X}(t)}{dt}$ is the system dynamics we want to learn. $T$ is the terminal time chosen as a hyperparameter. Specifically, $\frac{d\mathbf{X}(t)}{dt}$ is governed by a nonlinear opinion dynamics
where $d,\alpha$ are hyperparameters and $u=\frac{d}{\alpha+3}$. $\mathbf{A}^{\mathrm{a}}$ and $\mathbf{A}^{\mathrm{o}}$ are learnable adjacency matrices based on self-attention mechanism (Vaswani, 2017).
The numerical integration method used to solve the ODE is selected based on the dataset and can be one of forward Euler, Runge-Kutta 4th order (RK4), or the Dormand-Prince (Dopri5) method.
The decoder, $\psi:\mathbb{R}^{N_{\mathrm{o}}}\to\mathbb{R}^{d_{\text{class}}}$, is represented by a fully connected neural network followed by a softmax function, mapping the terminal opinion $\mathbf{X}(T)$ into the prediction $\mathbf{Y}$. For semi-supervised node classification task, we optimize the cross entropy loss
$\mathcal{L(\mathbf{Y,T})}=\sum_{i=1}^{n}\mathbf{t}_{i}^{\top}\text{log}\mathbf{y}_{i},$ (112)
where $\mathbf{t}_{i}\in\mathbb{R}^{d_{\text{class}}}$ is the one-hot truth vector of $i^{\text{th}}$ node.
C.2 Detail for Experiment 6.1
C.2.1 Detail for Dirichlet energy experiment
Experiment 6.1 illustrates the dynamics of Dirichlet energy in various GNNs, which indicates the similarity between the learned features. We randomly generate an undirected graph with 10 nodes represented by 2-dimensional features sampled from $\mathcal{U}[0,1]$. Also, we randomly initialize the models with the same seed and frozen them. The node features are then propagated forward through 1000 layers, and we plot the Dirichlet energy of each layer’s output with respect to the layer number, shown in Figure 2.
C.2.2 Detail for node classification across different depths
@@
Experiment 6.1 also evaluates various continuous-depth methods across different depths of $T=\{1,2,4,8,16,32,64,128\}$, while using the classification accuracy to measure the robustness to deep layers. Since adaptive step-size methods like Dormand–Prince (Dopri5) can result in varying numbers of integration steps, we use the Euler method with fixed step size $\Delta t=1$ for fair comparison.
However, the code implementation provided by $-\ell$ only supports Dormand–Prince (Dopri5), which may reduce its integration. While this introduces a potential unfairness, our method outperforms the baselines when the model reaches significant depths.
@@
For each method, we fix the set of hyperparameters across all depths. Without mentioning clearly, we use the fine-tuned parameters provided by each baseline. In BIMP, we use $d=0.5$ for Cora, $d=0.6$ for Citeseer and $b=0.2$ for Pubmed.
For all experiments, we run 100 train/valid/test splits for each dataset with 10 random seeds for each split on work stations with an Intel Xeon Gold 5220R 24 core CPU, an Nvidia A6000 GPUs, and 256GB of RAM.
C.3 Fine-tuned Parameters in Experiment 6.2
We search the hyperparameters using Ray Tune (Liaw et al., 2018) with 1000 random trails for each dataset. The fine tuned values for the damping term $d$ and final time $T$ are presented in Table 5.
Table 5: Fine-tuned key parameters. The damping term $d$ and time $T$ of 6 datasets used in Experiment 6.2.
Dataset Cora Citeseer Pubmed CoauthorCS Computers Photo
Damping Term $d$ 0.8952 1.0970 0.6908 0.1925 1.0269 1.0230
Time $T$ 12.2695 6.6067 9.7257 4.0393 3.2490 2.0281
C.4 Detail for Experiment 6.4
Noticing BIMP increase its expressive capacity by introducing an additional belief adjacency matrix $A^{o}$, its space complexity is higher than other baselines. However, given that the number of options is generally smaller than the number of agents, $A^{o}$ remains relatively small compared to the communication adjacency matrix $A^{a}$, resulting acceptable computational overhead.
Additionally, BIMP introduces nonlinearity through the saturation function $\text{tanh}(\cdot)$, which increases computational cost. However, considering this function operates element-wise, it is more efficient than other nonlinear dynamical processes, such as KuramotoGNN. In summary, although BIMP contains an belief matrix and a nonlinear function, it does not substantially increase computational complexity.
In order to illustrate the comparable computational complexity of BIMP, we record the running time for BIMP and other 4 popular continuous-depth GNNs, $-\ell$, $-\ell$, GraphCON-Tran and KuramotoGNN on Cora and Citeseer dataset. We firstly fine-tune the hyperparamters for each methods, than run each of them for 100 times with the fixed $epochs=100$ for Cora and $epochs=250$ for Citeseer. This experiment is deployed on work stations with an Intel Xeon Gold 5220R 24 core CPU, an Nvidia A6000 GPUs, and 256GB of RAM. The average running time listed in Table 6 demonstrates, in contrast to the nonlinear method KuramotoGNN, BIMP maintains a training time comparable to other linear continuous-depth baselines.

### v2

@@ C.1 Homophilic datasets
Table 2 in Section 6.1 performs semi-supervised node classification task on the Cora [85], citeseer [86], Pubmed [87], CoauthorCS [88] and Amazon Computers and Photo [93] homophilic datasets. While recent continuous-depth models such as $-\ell$ and KuramotoGNN makes use of the largest connected component of the datasets, we implement classification on full datasets.
@@
Cora. The Cora dataset contains a citation graph where 2708 computer science publications are connected by 5278 citation edges. Each publication has an 1433-dimensional bag-of-words vector derived from a paper keyword dictionary. Publications are classified into one of 7 classes corresponding to their primary research area.
@@
citeseer. The citeseer dataset contains a citation graph where 3312 computer science publications are connected by 4552 citation edges. Each publication has a 3703-dimensional bag-of-words vector derived from a paper keyword dictionary. Publications are classified into one of 6 classes corresponding to their primary research area.
@@
Pubmed. The Pubmed dataset contains a citation graph where 19717 biomedical publications are connected by 44324 citation edges. Each publication is represented by a 500-dimensional TF/IDF weighted word vector derived from a paper keyword dictionary. Publications are classified into one of 3 classes corresponding to their primary research area.
@@
CoauthorCS. The CoauthorCS dataset is one segment of the Coauthor Graph datasets that contains a co-authorship graph that consist of 18333 authors and connected by 81894 co-authorship edges. Each author is represented by a 6805-dimensional bag-of-words feature vector derived from their paper keywords. Authors are classified into one of 15 classes corresponding to their primary research area.
@@
Amazon Computers. The Amazon Computers dataset, denoted as Computers in our paper, contains a co-purchase graph where 13381 computer products are connected by 81894 edges. The edges indicate that two products are frequently bought. Each product is represented by a 767-dimensional bag-of-words feature vector derived from their product reviews. Products are classified into one of 10 classes corresponding to their product categories.
@@
Amazon Photo. The Amazon Photo dataset, denoted as Photo in our paper, contains a co-purchase graph where 7487 photo products are connected by 119043 edges. The edges indicate that two products are frequently bought. Each product is represented by a 745-dimensional bag-of-words feature vector derived from their product reviews. Products are classified into one of 8 classes corresponding to their product categories.
C.2 Heterophilic datasets
Table 3 in Section 6.1 performs semi-supervised node classification task on the Texas, Wisconsin, and Cornell heterophilic datasets from the CMU WebKB [89] project.
Texas. The Texas dataset contains a webpage graph where 183 web pages are connected by 325 hyperlink edges. Each webpage has a 1703-dimensional bag-of-words vector derived from the contents of the webpage. Webpages are classified into one of 5 classes corresponding to their primary content.
Wisconsin. The Wisconsin dataset contains a webpage graph where 251 web pages are connected by 512 hyperlink edges. Each webpage has a 1703-dimensional bag-of-words vector derived from the contents of the webpage. Webpages are classified into one of 5 classes corresponding to their primary content.
Cornell. The Cornell dataset contains a webpage graph where 183 web pages are connected by 298 hyperlink edges. Each webpage has a 1703-dimensional bag-of-words vector derived from the contents of the webpage. Webpages are classified into one of 5 classes corresponding to their primary content.
C.3 Large graphs
Table 6 in Appendix D.1.3 performs semi-supervised node classification task on the ogbn-arXiv [90] dataset. The ogbn-arXiv dataset consists of a single graph with 169,343 nodes and 1,166,243 edges where each node represents an arxiv paper, and edges represent citation relationships. We train each model in a semi-supervised way, and compute the training loss over 90,941 of the 169,343 nodes. We use 29,799 of the remaining nodes for validation, and the final 48,603 nodes for testing.
Appendix D Experiment Details
D.1 Classification accuracy
D.1.1 Classification accuracy on homophilic dataset
We demonstrate improved classification performance compared to $-\ell$, GRAND+++++ $-\ell$, KuramotoGNN, GraphCON, GAT, GCN and GraphSAGE. Additionally, we consider oversmoothing mitigation techniques of pairnorm [15] (denoted -pairnorm) and differentiable group normalization [16] (denoted -group) and training methods of adaptive attention (denoted -aa) and rewiring with graph diffusion convolution [94] (denoted -rw). Classification performance is reported in Table 4 where BIMP outperforms baseline methods.
Table 4: Classification accuracy on homophilic datasets. Classification accuracies on the Cora, citeseer, Pubmed, CoauthorCS, Computers, and Photo datasets are reported, where BIMP outperforms competitive baselines.
Dataset Cora citeseer Pubmed CoauthorCS Computers Photo
BIMP 83.19$\pm$1.13 71.09$\pm$1.40 80.16$\pm$2.03 92.48$\pm$0.26 84.73$\pm$0.61 92.90$\pm$0.44
BIMP-aa 82.96$\pm$1.31 70.43$\pm$0.80 80.35$\pm$0.99 91.82$\pm$0.37 84.72$\pm$0.45 92.27$\pm$0.36
BIMP-aa-rw 82.59$\pm$1.06 70.51$\pm$1.37 78.56$\pm$1.12 91.97$\pm$0.37 84.76$\pm$0.23 92.92$\pm$0.19
$-\ell$ 82.20$\pm$1.45 69.89$\pm$1.48 78.19$\pm$1.88 90.23$\pm$0.91 82.93$\pm$0.56 91.93$\pm$0.39
GRAND-aa 82.59$\pm$0.28 70.21$\pm$1.21 78.39$\pm$1.95 91.44$\pm$0.42 83.09$\pm$1.71 92.50$\pm$0.53
GRAND-aa-rw 82.86$\pm$1.47 70.95$\pm$1.13 78.56$\pm$1.13 91.52$\pm$0.31 83.47$\pm$0.51 92.64$\pm$0.24
$-\ell$ 82.83$\pm$1.31 70.26$\pm$1.46 78.89$\pm$1.96 90.10$\pm$0.78 82.79$\pm$0.54 91.51$\pm$0.41
GRAND++-aa 80.14$\pm$0.93 69.94$\pm$1.45 78.50$\pm$1.28 85.65$\pm$1.30 84.00$\pm$0.47 91.86$\pm$0.52
GRAND++-aa-rw 81.91$\pm$1.39 69.41$\pm$0.95 79.44$\pm$1.06 86.23$\pm$0.80 83.35$\pm$0.63 92.50$\pm$0.22
KuramotoGNN 81.16$\pm$1.61 70.40$\pm$1.02 78.69$\pm$1.91 91.05$\pm$0.56 80.06$\pm$1.60 92.77$\pm$0.42
GraphCON-Tran 82.80$\pm$1.34 69.60$\pm$1.16 78.85$\pm$1.53 90.30$\pm$0.74 82.76$\pm$0.58 91.78$\pm$0.50
GAT 79.76$\pm$1.50 67.70$\pm$1.63 76.88$\pm$2.08 89.51$\pm$0.54 81.73$\pm$1.89 89.12$\pm$1.60
GCN 80.76$\pm$2.04 67.54$\pm$1.98 77.04$\pm$1.78 90.98$\pm$0.42 82.02$\pm$1.87 90.37$\pm$1.38
GCN-pairnorm 79.55$\pm$1.21 66.93$\pm$0.94 76.14$\pm$0.63 90.63$\pm$0.69 81.88$\pm$2.73 86.93$\pm$1.35
GCN-group 80.48$\pm$1.40 66.99$\pm$1.97 77.53$\pm$0.97 90.97$\pm$0.54 81.97$\pm$0.75 89.84$\pm$0.71
GraphSAGE 79.37$\pm$1.70 67.31$\pm$1.63 75.52$\pm$2.19 90.62$\pm$0.42 76.42$\pm$7.60 88.71$\pm$2.68
D.1.2 Classification accuracy on heterophilic dataset
We demonstrate improved classification performance compared to $-\ell$, GRAND+++++ $-\ell$, KuramotoGNN, GraphCON, GAT, GCN and GraphSAGE. Additionally, we consider oversmoothing mitigation techniques of pairnorm (denoted -pairnorm) and differentiable group normalization (denoted -group) and training methods of adaptive attention (denoted -aa) and rewiring with graph diffusion convolution (denoted -rw). Classification performance is reported in Table 5 where BIMP outperforms baseline methods.
Table 5: Classification accuracy on heterophilic datasets. Classification accuracies on the Texas, Wisconsin, and Cornell datasets are reported, where BIMP outperforms competitive baselines.
Dataset Cornell Texas Wisconsin
Homophily level 0.30 0.11 0.21
BIMP 77.13$\pm$3.38 82.16$\pm$4.06 86.57$\pm$4.33
BIMP-aa 76.95$\pm$4.71 82.25$\pm$6.49 86.27$\pm$4.36
BIMP-aa-rw 77.46$\pm$4.80 82.43$\pm$6.76 86.22$\pm$4.34
$-\ell$ 70.00$\pm$6.22 74.59$\pm$5.43 82.75$\pm$3.90
GRAND-aa 71.89$\pm$5.30 75.68$\pm$9.89 82.16$\pm$3.77
GRAND-aa-rw 74.59$\pm$4.22 82.16$\pm$5.43 83.73$\pm$5.69
$-\ell$ 70.30$\pm$8.50 76.14$\pm$5.77 83.09$\pm$2.83
GRAND++-aa 71.89$\pm$3.43 78.57$\pm$6.51 83.96$\pm$5.02
GRAND++-aa-rw 74.05$\pm$5.57 79.42$\pm$5.24 84.62$\pm$3.56
KuramotoGNN 76.02 $\pm$2.77 81.81$\pm$4.36 85.09$\pm$4.42
GraphCON-GCN 74.05$\pm$3.24 80.54$\pm$4.49 84.79$\pm$2.51
GAT 42.16$\pm$7.07 57.84$\pm$5.82 49.61$\pm$4.21
GCN 41.35$\pm$4.69 57.03$\pm$5.98 48.43$\pm$5.75
GCN-pairnorm 52.70$\pm$6.42 63.51$\pm$6.54 60.59$\pm$4.34
GCN-group 47.62$\pm$5.30 59.92$\pm$4.15 51.37$\pm$5.67
GraphSAGE 70.54$\pm$2.55 72.70$\pm$5.47 73.14$\pm$6.27
D.1.3 Experiment on large graph
We demonstrate improved classification performance compared to $-\ell$, KuramotoGNN, GCN and GAT on large graphs. Classification performance is reported in Table 6 where BIMP outperforms baseline methods.
Table 6: Classification accuracy on ogbn-arXiv dataset. Our BIMP model outperforms $-\ell$, KuramotoGNN, GCN and GAT on the ogbn-arXiv dataset.
Dataset ogbn-arXiv
BIMP 71.04$\pm$0.94
$-\ell$ 70.19$\pm$0.43
KuramotoGNN 66.96$\pm$0.25
GCN 61.66$\pm$0.32
GAT 69.86$\pm$0.59
D.1.4 Homophilic dataset hyperparameters
We search the hyperparameters using Ray Tune [95] with 1000 random trials for each dataset and final values are shown in Table 7. Experiments were run with 100 random splits and each split trained on 100 seeds (seed value for both random split and training are 1-100).
Table 7: Hyperparameter for homophilic dataset. The hyperparameters for homophilic datasets in Section 6.1 is reported.
Dataset Cora citeseer Pubmed CoauthorCS Computers Photo
Opinion Dim. $80$ $128$ $128$ $16$ $128$ $64$
Epoch $100$ $250$ $600$ $250$ $100$ $100$
Learning Rate $0.0178$ $0.0034$ $0.0210$ $0.0018$ $0.0035$ $0.0056$
Optimizer AdaMax AdaMax AdaMax RMSProp Adam Adam
Weight Decay $0.0078$ $0.1$ $0.0020$ $0.0047$ $0.0077$ $0.0047$
Dropout $0.1353$ $0.3339$ $0.0932$ $0.6858$ $0.0873$ $0.4650$
Input Dropout $0.4172$ $0.5586$ $0.6106$ $0.5275$ $0.5973$ $0.4290$
Attention Head $4$ $2$ $1$ $4$ $4$ $4$
Attention Dim. $16$ $8$ $16$ $8$ $64$ $64$
Attention Type Scaled Dot Exp. Kernel Cosine Sim. Scaled Dot Scaled Dot Pearson
NODE Adjoint False False True True True True
Adjoint Method n/a n/a Euler dopri5 dopri5 dopri5
Adjoint Step Size n/a n/a $1$ $1$ $1$ $1$
Integral Method dopri5 dopri5 dopri5 dopri5 dopri5 dopri5
Linear Encoder True True False True True False
Linear Decoder True True True True True True
Step Size $1$ $1$ $1$ $1$ $1$ $1$
Time ($T$) $12.2695$ $6.6067$ $9.7257$ $4.0393$ $3.2490$ $2.0281$
Damping ($d$) $0.8952$ $1.0970$ $0.6908$ $0.1925$ $1.0269$ $1.0230$
Self-reinforce ($\alpha$) $1$ $1$ $1$ $1$ $1$ $1$
D.1.5 Heterophilic dataset hyperparameters
We search the hyperparameters using Ray Tune with 200 random trials for each dataset and final values are shown in Table 8. Experiments were run with 10 standardized splits and each split trained on 100 seeds (seed value for training are 1-100).
Table 8: Hyperparameter for heterophilic dataset. The hyperparameters for heterophilic datasets in Section 6.1 is reported.
Dataset Texas Wisconsin Cornell
Opinion Dim. $256$ $32$ $32$
Epoch $200$ $100$ $100$
Learning Rate $0.0178$ $0.0178$ $0.0218$
Optimizer AdaMax AdaMax AdaMax
Weight Decay $0.0078$ $0.0091$ $0.0478$
Dropout $0.6531$ $0.2528$ $0.2030$
Input Dropout $0.0052$ $0.0042$ $0.0417$
Attention Head $8$ $4$ $4$
Attention Dim. $32$ $16$ $16$
Attention Type Scaled Dot Scaled Dot Scaled Dot
NODE Adjoint False False False
Integral Method dopri5 dopri5 dopri5
Linear Encoder False False False
Linear Decoder True False False
Step Size $1$ $1$ $1$
Time ($T$) $0.01$ $0.01$ $0.01$
Damping ($d$) $0.0086$ $0.0075$ $0.0195$
Self-reinforce ($\alpha$) $2$ $2$ $1.5$
D.2 Performance at large depths
D.2.1 Classification accuracy
Experiment 6.2: Classification accuracy evaluates the classification performance of BIMP and continuous-depth baselines at different depths of $T=\{1,2,4,8,16,32,64,128\}$ with 100 splits and 10 random seeds. We use the classification accuracy as a measure of the robustness to deep layers of BIMP and baseline methods.
Figure 2 show the comparison of classification accuracy of BIMP and select continuous baseline methods. Figure 3 show the comparison of classification accuracy of BIMP and additional baselines and oversmoothing mitigation techniques.
@@ D.2.1 Classification accuracy
Since adaptive step-size methods like Dormand–Prince (Dopri5) can result in inconsistent numbers of integration steps, we implement the Euler method with fixed step size $\Delta t=1$ for fair comparison. Notably, BIMP outperforms $-\ell$ at significant depths, even though $-\ell$ only supports Dormand–Prince (Dopri5).
@@
For each method, we use the fine-tuned parameters provided by each baseline and fix the set of hyperparameters across all depths. For all experiments, we run 100 train/valid/test splits for each dataset with 10 random seeds for each split.
Figure 3: Classification accuracy. BIMP is designed to learn node representations that resist oversmoothing even for very large depths. We compare the classification accuracy of BIMP to baseline models for architectures with $1,2,4,8,16,32,64$ and $128$ timesteps. Our BIMP model and its variants are stable out to 128 timesteps, while baseline performance deteriorates after 32 timesteps.
Figure 4: Dirichlet energy. BIMP is designed to learn node representations that resist oversmoothing even for very large depths. We compare the Dirichlet energy of node features over a range of network depths. The Dirichlet energy of BIMP remains stable even at very deep layers, while the energy of baseline modes does not.
D.2.2 Dirichlet energy
Experiment 6.2: Dirichlet energy illustrates the dynamics of the Dirichlet energy in BIMP and baseline methods, which indicates the similarity between the learned features. We randomly generate an undirected graph with 10 nodes each with 2-dimensional features sampled from $\mathcal{U}[0,1]$. We randomly initialize the models with the same seed and the node features are propagated forward through 1000 layers.
Figure 2 show the comparison of Dirichlet energy of BIMP and select continuous baseline methods. Figure 4 show the comparison of Dirichlet energy of BIMP and additional baselines and oversmoothing mitigation techniques.
Appendix E Additional Experiments
E.1 Computational complexity
BIMP’s space complexity is higher than other baselines as its increase in expressive capacity comes from the introducing of an additional option adjacency matrix $\mathbf{A}^{\mathrm{o}}$. Specially, we compare the space complexity of BIMP with that of $-\ell$. The complexity of BIMP is $\mathcal{O}(mN_{\mathrm{o}}+N_{\mathrm{o}}^{2})$, where $m$ is the number of edges, $N_{\mathrm{a}}$ is the number of nodes, and $N_{\mathrm{o}}$ is the number of options. The space complexity of $-\ell$ is $\mathcal{O}(mN_{o})$. Given that the number of options is generally smaller than the number of agents, $N_{o}$ remains relatively small compared to the number of edge $m$, resulting acceptable computational overhead.
Additionally, BIMP introduces a nonlinearity through the saturation function $\text{tanh}(\cdot)$, which increases computational cost. However, considering this function operates element-wise, it is more efficient than other nonlinear dynamical processes, such as KuramotoGNN. To illustrate this difference, we compare the average run time of our model against baseline models in Table 9. We note that BIMP has comparable run time performance to linear baselines models such as $-\ell$.
We record the running time for BIMP and other 4 popular continuous-depth GNNs, $-\ell$, $-\ell$, GraphCON-Tran and KuramotoGNN on Cora and citeseer dataset. We train each model 100 times with with a fixed number of epoch (100 for Cora and 250 for citeseer) using fine tuned hyperparameters. The average training time listed in Table 9 demonstrates, in contrast to other nonlinear methods like KuramotoGNN, BIMP maintains a training time comparable to other linear continuous-depth baselines.
All experiments reported in the paper was conducted on work stations with an Intel Xeon Gold 5220R 24 core CPU, an Nvidia A6000 GPUs, and 256GB of RAM.